Shift report 20011122 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Nov 22 2001 - 07:52:05 EST

  • Next message: Jens Jørgen Gaardhøje: "width"

    ------------------------------------------------------------
    Date: Thu Nov 22 07:52:04 2001, Shift: 00:00-08:00
    Supervisor: Andrei Makeev, with: 
    Log book#11 pages 23 - 23
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 0:20 a.m. a cryogenic problem has occured in yellow
    ring. Beam died, run #5938 was stopped. At 4:12 a.m.
    some troubles appeared with injection into AGS. Repair
    took remaining 4 hours... no beam for whole shift long.
    2. No achievements.
    3. No detector problems.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Nov 22 2001 - 07:52:30 EST